A coin bank contains only nickles and dimes the # of dims is 2 less than twice the # of nickles. There are 52 coins in the bank How money is in the bank
Found 2 solutions by josmiceli, funmath:
Answer by josmiceli(19441)

d + n = 52
d = 2n - 2
d = 52 - n
52 - n = 2n - 2
3n = 54
n = 18
d = 34
check
d = 2n - 2
34 = 2*18 - 2
34 = 34
OK

Answer by funmath(2933)

The number of nickles is: N
The number of dimes is 2 less than twice the nickles: 2N-2
(The number of nickles)+(the number of dimes)=52 Therefore:
N+2N-2=52
3N-2=52
3N-2+2=52+2
3N=54
3N/3=54/3
N=18
The number of nickles "N" is 18
The number of dimes is 2N-2
Substitute 18 in for N and you have: 2(18)-2=34
Nickles are worth 5 cents, so you have 5(18)=90 cents in nickels
Dimes are worth 10 cents, so you have 10(34)=340 cents in dimes
90+340=430 cents all together or $4.30